Calculate the volume of a cuboidal box whose dimensions are 5x × 3x2 × 7x4

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Length = 5x

Breadth = 3x^2

Height = 7x^4

Area of cuboid = Length × Breadth × Height

= 5x × 3x^2× 7x^4

Multiply 5, 3, and 7

= 105 x x^2 x^4

Use exponents rule: x^a× x^b= x^(a+b)

So, 105 x x^2 x^4= 105x^(1+2+4) = 105x^7

Thanks